Circular No. 7266

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V4334 SAGITTARII K. Hinkle and R. Joyce, Kitt Peak National Observatory (KPNO), re-

port infrared photometry of V4334 Sgr (cf. IAUC 6322) obtained on two different instruments (TIFKAM on Sept. 1, observers Joyce and A. Koekemoer; and the Phoenix spectrograph in imaging mode with narrow-band filters on the other nights, observers Hinkle, C. Kulesa, and J. Melendez) at the KPNO 2.1-m telescope: Sept. 1.15 UT, J = 13.3, H = 9.9, K = 7.2; 18.09, [2230 nm] = 6.9; 19.10, [1558 nm] = 10.5; 19.11, [1282 nm] = 13.0; 27.09, [3226 nm] = 4.2; 28.08, [4651 nm] = 2.5. The photometry has an accuracy of better than 0.2 mag. V4334 Sgr is fainter at K than in April (cf. IAUC 7155) and has become redder. The infrared photometry is consistent with a 740-K blackbody.

GRS 1758–258 D. M. Smith, University of California, Berkeley; W. A. Heindl, Uni-

versity of California, San Diego; and J. H. Swank, Goddard Space Flight Center, write: “Weekly RXTE PCA monitoring of blackhole candidate GRS 1758–258 shows that this persistent source, which has been in a hard state since at least the beginning of 1996, has since 1999 Feb. 19 entered a period of oscillation between hard (power-law index 1.4–1.7, thermal disk component undetectable or < 10 percent of the 2–10-keV luminosity) and soft (index 2.0–2.3, thermal disk > 20 percent) states. The first cycles had only minor spectral changes and a 1-month timescale. The changes have become gradually more extreme, and, since Sept. 12, have begun to occur on scales < 1 week. The state changes may modulate mass loss to the radio jets. Observations in the x-ray and infrared-to-radio bands are strongly encouraged, frequent monitoring in particular.”

COMET C/1999 H1 (LEE) Extension to ephemeris on IAUC 7251 (elements MPC 35814):

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 29 1^{h}29.^{m}80 +55^{o}19.^{′}3 0.828 1.625 125.^{o}0 30.^{o}3 8.7
Oct. 4 0 32.25 +49 27.6 0.837 1.697 134.6 24.8 8.9
9 23 50.62 +42 22.5 0.877 1.768 140.8 20.9 9.2
14 23 21.83 +35 16.0 0.944 1.838 142.4 19.3 9.5
19 23 02.06 +28 50.9 1.036 1.908 139.8 19.7 9.9
1999 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green